# Structuring the Unstructured: A Multi-Agent System for Extracting and Querying Financial KPIs and Guidance

### v1

@@ Abstract.
Extracting structured and quantitative insights from unstructured financial filings is essential in investment research, yet remains time-consuming and resource-intensive. Conventional approaches in practice rely heavily on labor-intensive manual processes, limiting scalability and delaying the research workflow. In this paper, we propose an efficient and scalable method for accurately extracting quantitative insights from unstructured financial documents, leveraging a multi-agent system composed of large language models. Our proposed multi-agent system consists of two specialized agents: the Extraction Agent and the Text-to-SQL Agent. The Extraction Agent automatically identifies key performance indicators from unstructured financial text, standardizes their formats, and verifies their accuracy. On the other hand, the Text-to-SQL Agent generates executable SQL statements from natural language queries, allowing users to access structured data accurately without requiring familiarity with the database schema. Through experiments, we demonstrate that our proposed system effectively transforms unstructured text into structured data accurately and enables precise retrieval of key information. First, we demonstrate that our system achieves approximately 95% accuracy in transforming financial filings into structured data, matching the performance level typically attained by human annotators. Second, in a human evaluation of the retrieval task – where natural language queries are used to search information from structured data – 91% of the responses were rated as correct by human evaluators. In both evaluations, our system generalizes well across financial document types, consistently delivering reliable performance.
@@ 1. Introduction
The financial industry generates vast volumes of unstructured documents-10-Ks, 10-Qs, 8-Ks, DEF14As, and investor memos-that are central to investment research and decision-making. Analysts have traditionally extracted key performance indicators (KPIs) from these documents by manually scanning text, copying metrics, normalizing formats, and writing internal summaries. This process is not only time-consuming but also error-prone and unscalable in modern research environments where timeliness and breadth of coverage are essential.
While recent large language models (LLMs) have shown promise in structuring unstructured data, their application in financial workflows remains unreliable. In the extraction phase, LLMs often misinterpret numeric guidance, extract semantically incorrect values (e.g., YoY instead of QoQ), or omit critical metrics entirely. These issues stem from a lack of domain-specific reasoning, layout ambiguity, and difficulty distinguishing factual signals from narrative commentary. Even when extraction is successful, querying this data via natural language introduces further challenges: text-to-SQL models are sensitive to phrasing, schema variation, and inherent LLM randomness, often leading to invalid or misleading responses.
These problems are compounded by the fact that most LLM-based systems rely on monolithic, single-shot inference pipelines that lack transparency, fail to isolate errors, and cannot iteratively refine intermediate outputs. This makes it difficult to enforce consistency, inject domain rules, or apply structured validation across stages.
To address these challenges, we propose a scalable, document-agnostic multi-agent framework for financial document understanding. Our system consists of two specialized agents: an Extraction Agent, which parses and validates KPIs using domain-tuned prompts and logic, and a Text-to-SQL Agent, which translates analyst queries into executable SQL over the structured output. By explicitly decomposing the workflow, the system allows targeted evaluation, modular tuning, and the integration of feedback loops at each stage-enabling iterative refinement and robust error handling.
This modular design not only automates manual workflows but also ensures both high-fidelity structuring and precise, trustworthy retrieval. Evaluated on a diverse set of SEC filings and real analyst tasks, our system demonstrates significant improvements in accuracy, controllability, and scalability over traditional and monolithic LLM baselines.
@@ 3.3. Extraction Agent
The Extraction Agent transforms raw financial documents into structured, machine-readable KPI records through a seven-stage process combining neural reasoning and rule-based logic. Fig. 3 shows performance across each step.
Figure 3. Evaluation breakdown for the Extraction Agent across its seven iterative stages. Each step contributes to higher accuracy, schema alignment, and unit/period consistency.
@@ 3.4. Text-to-SQL Agent
The Text-to-SQL Agent enables analysts to pose natural language questions over the structured KPI database. It translates queries like “What was the Q4 operating margin?” into executable SQL, handling financial semantics and schema variation.
@@
Figure 4. Iterative pipeline of the Text-to-SQL Agent. Each stage—from intent parsing and schema enrichment to constraint validation and execution testing—is designed to ensure high-fidelity SQL generation and interpretability.
@@ 4.1. Experimental Setup
We compare three configurations:
@@ 4.2. Component-Level Evaluation and End-to-End Pipeline Comparison
Config. Extraction F1 Query Accuracy Unit Error Rate Throughput Gain
GPT-4 (naive zero-shot) 72.4% 81.3% 18.5% 1.0×
Claude 3.7 (domain prompt) 84.7% 89.6% 7.2% 1.5×
Claude + Rule Injection 91.2% 94.8% 2.3% 2.4×
We first assess each agent in isolation to measure its contribution to overall accuracy in Table 1 and  2. Next, we evaluate the combined system across the three configurations in Table 3. The results show that structured prompting and rule-based validation significantly improve both extraction and query accuracy while reducing error rates and increasing throughput.
Prompt Engineering Impact.
Rule Injection Effect.
Query Reliability.
System Efficiency.

### v2

@@ Abstract.
Extracting and retrieving structured and quantitative insights from unstructured financial filings is essential in investment research, yet remains time-consuming and resource-intensive. Conventional approaches in practice rely heavily on labor-intensive manual processes, limiting scalability and delaying the research workflow. In this paper, we propose an efficient and scalable method for both extracting and retrieving financial insights accurately, leveraging a multi-agent system composed of large language models. Our proposed multi-agent system consists of two specialized agents: the Extraction Agent and the Text-to-SQL Agent. The Extraction Agent automatically identifies key performance indicators from unstructured financial text, standardizes their formats, and verifies their accuracy. On the other hand, the Text-to-SQL Agent generates executable SQL statements from natural language queries, allowing users to access structured data accurately without requiring familiarity with the database schema. Through experiments, we demonstrate that our proposed system effectively transforms unstructured text into structured data accurately and enables precise retrieval of key information. First, we demonstrate that our system achieves approximately 95% accuracy in transforming financial filings into structured data, matching the performance level typically attained by human annotators. Second, in a human evaluation of the retrieval task – where natural language queries are used to search information from structured data – 91% of the responses were rated as correct by human evaluators. In both evaluations, our system generalizes well across financial document types, consistently delivering reliable performance.
@@ 1. Introduction
Investment research depends critically on the analysis of unstructured corporate disclosures. In practice, analysts often extract key performance indicators (KPIs) either by manually scanning lengthy documents or by locating figures scattered across external sources and internal summaries. As firms continue to release a growing volume of earnings transcripts, financial reports, and regulatory filings, such manual processes have become increasingly time-consuming and impractical. To ensure both timeliness and breadth of coverage, scalable methods for extracting and retrieving insights from these texts are now essential to informed decision-making. Recent advances in large language models (LLMs) have opened new possibilities for processing unstructured documents at scale, offering a promising foundation for automating financial information extraction and retrieval. These models can, in principle, transform complex textual data into structured insights and support natural language interactions with document content.
However, despite their potential, LLMs yet remain unreliable in real-world financial workflows. When used in isolation, they often misinterpret numeric guidance, extract semantically incorrect values, or omit critical metrics altogether. Such failures are largely due to limitations in layout sensitivity and the difficulty of distinguishing factual signals from surrounding narrative. To address these shortcomings, recent work has explored multi-agent frameworks, in which multiple LLMs collaborate to validate, cross-check, and refine each other’s outputs—improving accuracy, reducing hallucinations, and enhancing reliability in financial document analysis.
In this context, we propose an efficient and scalable method for both extracting and retrieving financial insights accurately, leveraging a multi-agent system composed of LLMs. Our system consists of two specialized agents: an Extraction Agent, which parses and validates KPIs using domain-tuned prompts and logic, and a Text-to-SQL Agent, which translates analyst queries into executable SQL over the structured output. By explicitly decomposing the workflow, the system allows targeted evaluation, modular tuning, and the integration of feedback loops at each stage-enabling iterative refinement and robust error handling. This modular design not only automates manual workflows but also ensures both high-fidelity structuring and precise, trustworthy retrieval. Evaluated on a diverse set of SEC filings and real analyst tasks, our system demonstrates significant improvements in accuracy, controllability, and scalability over traditional and monolithic LLM baselines.
@@ 3.2. System Overview: A Two-Agent Architecture
Figure 3. Overview of our multi-agent system for financial document understanding. The Extraction Agent processes unstructured documents to extract structured KPIs using LLMs (LLaMA, Claude) and stores them in Delta Lake via Apache Spark. The Text-to-SQL Agent enables natural language querying by generating metadata-aware SQL queries, retrieving results through Databricks SQL. The system is built on the Databricks stack, including MLflow for auditing and AI Gateway for traceability.
@@ 3.3. Extraction Agent
The Extraction Agent transforms raw financial documents into structured, machine-readable KPI records through a seven-stage process combining neural reasoning and rule-based logic. The top side of Fig. 3 shows its detailed steps.
@@ 3.4. Text-to-SQL Agent
The Text-to-SQL Agent enables analysts to pose natural language questions over the structured KPI database. It translates queries like “What was the Q4 operating margin?” into executable SQL, handling financial semantics and schema variation. The bottom side of Fig. 3 shows its detailed steps.
@@ 4. Experiments
Config. Extraction F1 Query Accuracy Unit Error Rate Throughput Gain
GPT-4o (naive zero-shot) 72.4% 81.3% 18.5% 1.0×
Claude-3.7-Sonnet (domain prompt) 84.7% 89.6% 7.2% 1.5×
Claude-3.7-Sonnet + Rule Injection 91.2% 94.8% 2.3% 2.4×
@@ 4.1. Experimental Setup
Each document was manually annotated by trained financial analysts to identify ground-truth KPIs, including revenue, operating income, free cash flow, and forward-looking guidance. Annotations were cross-validated for consistency and aligned with GAAP definitions where applicable. We compare three configurations:
@@ 4.2. Component-Level Evaluation and End-to-End Pipeline Comparison
We first assess each agent in isolation to measure its contribution to overall accuracy in Table 2 and  3. Next, we evaluate the combined system across the three configurations in Table 1. The results show that structured prompting and rule-based validation significantly improve both extraction and query accuracy while reducing error rates and increasing throughput.
Prompt Engineering Impact.
Rule Injection Effect.
Query Reliability.
System Efficiency.